# Alternative Approaches for Computing Highest-Density Regions

### v1

@@ Abstract
Many statistical problems require estimating a density function, say $f$, from data samples. In this work, for example, we are interested in highest-density regions (HDRs), i.e., minimum volume sets that contain a given probability. HDRs are typically computed using a density quantile approach, which, in the case of unknown densities, involves their estimation. This task turns out to be far from trivial, especially over increased dimensions and when data are sparse and exhibit complex structures (e.g., multimodalities or particular dependencies). We address this challenge by exploring alternative approaches to build HDRs that overcome direct (multivariate) density estimation. First, we generalize the density quantile method–currently implementable on the basis of a consistent estimator of the density–to neighbourhood measures, i.e., measures that preserve the order induced in the sample by $f$. Second, we discuss a number of suitable probabilistic- and distance-based measures such as the $k$-nearest neighbourhood Euclidean distance. Third, motivated by the ubiquitous role of copula modeling in modern statistics, we explore its use in the context of probabilistic-based measures. An extensive comparison among the introduced measures is provided, and their implications for computing HDRs in real-world problems are discussed.
Keywords Statistical intervals  $\cdot$ Density estimation  $\cdot$ Copula models  $\cdot$ k-nearest neighborhood
@@ 1 Introduction
A ubiquitous problem in statistics is to derive statistical intervals or regions–especially in the multivariate setting–for population parameters or other unknown quantities. Their role is to provide a way to quantify and describe the uncertainty about a quantity of interest, or simply a way to summarize the information contained in a distribution. Statistical regions may address different problems. For example, a confidence interval (CI) describes the uncertainty surrounding an estimate for an unknown parameter, while a prediction interval provides bounds for one or more future observations. Alternatively, a tolerance interval is the interval expected to contain a specified proportion of the sampled population. In a Bayesian setting, highest posterior density or credible regions provide, in a natural way, set estimates for a specific parameter (Box and Tiao,, 1992; Turkkan and Pham-Gia,, 1993). We refer to Meeker et al., (2017) and Krishnamoorthy and Mathew, (2009) for an overview. Furthermore, even when the focus is on a specific type of interval, e.g., a two-sided 95% prediction interval, questions on the way this should be defined may still arise. Should we use the interval symmetric about the mean, the interval of shortest length, or the interval that minimizes the probability of covering a given set? Although each of these intervals has 95% coverage, they may all be different. Consider, for example, the case of nonsymmetric or/and multimodal distributions, with an illustration given in Figure 1. An additional layer of complexity arises in multivariate settings, where there are no unique agreed definitions, and generalizations include the concept of simultaneous CIs (Guilbaud,, 2008), multivariate CIs (Korpela et al.,, 2017), or different definitions for multivariate quantiles (Figalli,, 2018; Cai,, 2010; Coblenz et al.,, 2018), among others.
@@
The multivariate setting will be the target of this work, with a focus on bivariate distributions. In particular, our interest is devoted to statistical regions for summarizing probability distributions in the form of highest-density regions (HDRs; Hyndman,, 1996). Statistical regions other than HDRs are beyond the scope of the present work. As the name suggests, an HDR specifies the set of points of highest density: The density for points inside the region must be higher than that for points outside it. More specifically, considering a $d$-dimensional continuous variable of interest $X\in\mathbb{R}^{d}$ with probability density function $f$, the concrete problem is to estimate minimum volume sets of the form $C(f_{\alpha})=\{x\!:f(x)\geq f_{\alpha}\}$, such that $P(X\in C(f_{\alpha}))\geq 1-\alpha$, where $1-\alpha$, with $\alpha\in(0,1)$, represents a prespecified coverage probability. Although they share substantial similarities with multivariate quantiles, estimating an HDR differs from estimating level sets in that one is interested in specifying a probability content rather than the level directly. This complicates the problem, and we refer to Doss and Weng, (2018) for more details.
@@
To detect abnormal observations from a sample: if a data point does not belong to a region of “normal” or concentrated data (the HDR), then it is regarded anomalous (see e.g., Steinwart et al.,, 2005, and references therein).
@@
Due to their flexibility “to convey both multimodality and asymmetry”, HDRs are argued to be a more effective summary of the distribution (Hyndman,, 1995). In the case of unimodal symmetric distributions, such as the normal distribution, an HDR coincides with the usual probability region symmetric about the mean, spanning the $\alpha/2$ and $21-\alpha/$ 2 quantiles. However, in the case of a multimodal distribution, it may consist of several disjoint subregions, each containing a local mode. This provides useful information that is “masked” by other probability regions. An illustrative example is provided in Figure 1, where an HDR is compared to an equal-tailed interval.
@@
In principle, knowledge of an accurate estimator of the underlying probability density function $f$ of a variable $X$ is sufficient to compute its HDRs. However, existing methods for estimating $f$, such as the kernel density estimator (KDE; Parzen,, 1962) or the local likelihood approach (Hjort and Jones,, 1996), work very well for unidimensional problems, but may be inefficient for multidimensional problems (Liu et al.,, 2007). In fact, over increased dimensions, KDE suffers from the difficulty of finding optimal kernel functions and the corresponding bandwidths (i.e., the smoothing parameters). In particular, bandwidth selection in KDE is recognized as the most crucial and difficult step (see e.g., Chapter 2 in Wand and Jones, 1994a, ), with no definite and broadly accepted solution. As an illustrative example, in Figure 2 we estimate a $95\%$ HDR for a bivariate Gaussian distribution using KDE and three different bandwidth specifications. These are given in the works of Silverman, (1996)–in the form of rule of thumb, in Wand and Jones, 1994b –currently employed in the ks and hdrcde packages in R, and in Chacón et al., (2011)–under optimal asymptotic considerations. As depicted in Figure 2, the three different values result in three different HDRs, with an open question on the most appropriate one.
Figure 2: Estimated 95% HDR on a sample of $n=200$ units using KDE with three bandwidth specifications. Cadet-blue points are the HDR points, while purple x-marks and red triangles denote the points outside the HDR. The blue contour line represents the true level set (i.e., the one delimiting a 95% density region) of a bivariate Gaussian distribution (see Scenario S2 in Section 4).
In particular, if the task would be to identify abnormal or anomalous values, contradicting conclusions would arise from the three different HDRs (see e.g., the red triangle points in Figure 2). More concretely, our analysis is directly motivated by, and sees our involvement in, the World Anti-Doping Agency (WADA) mission of doping detection.
Furthermore, high-dimensional data also pose challenges from an algorithmic/computational perspective. Taken together, these aspects hamper the ability to derive an appropriate HDR. Motivated by these existing density-estimation issues, our research question is the following: “Can we relax the multivariate density estimation problem, either by using alternative ways (to direct density estimation, e.g., using distance-based proxies) that will robustly inform the density around a given point or by reducing the multidimensional estimation problem to a lower dimension?”. We address this question by exploring, evaluating, and proposing some novel approaches to building HDRs.
@@
The contribution of the present work is fourfold. First, we generalize the density-quantile approach for computing HDRs–currently implementable on the basis of a consistent estimator of the density–to neighbourhood measures (Munoz and Moguerza,, 2006), i.e., measures that preserve the order induced in the sample by the density function (Section 2). Second, we elaborate on and evaluate a number of suitable probabilistic- and distance-based measures, including a variation of the measure adopted by the k-nearest neighbors algorithm (Section 3). Third, motivated by the ubiquitous role of copula modeling (Nelsen,, 2006) in modern statistics, among probabilistic-based measures, we explore the use of copulae in an HDR estimation context. This is done in Section 3.4. Specifically, copulae make it possible to relax the estimation of multivariate random vectors, by separately estimating the marginals and their dependence structure, i.e., the copula model. In addition, by placing a strong focus on the dependence model, a copula approach has the advantage of better capturing data specificities, especially when these exhibit complex relationships such as asymmetric and/or tail dependencies. Finally, in Section 4, we provide a comprehensive comparison between the introduced measures and standard density estimators in all simulated scenarios, as well as in the MAGIC data set, which classifies high-energy Gamma particles in the atmosphere (Bock et al.,, 2004). Empirical studies are focused on bivariate scenarios that vary according to the complexity of the data (marginal and dependence structure, multimodality, etc) and the sample size. A distribution on a compact support, namely, the Dirichlet distribution, is also considered. We conclude by summarizing the main findings and discussing their implications for computing HDRs in real-world problems.
2 HDR: Formulation and Computation
@@
Let $\{X_{1},\dots,X_{n}\}$ be a sample of $n$ independent and identically distributed (iid) replications of a random variable $X$ defined on $\mathbb{R}^{d}$, with $d\geq 1$. In our problem, we assume to have access to a sample $s_{n}=\{x_{1},\dots,x_{n}\}\in S_{n}$ of their actual realizations, which we wish to use to build an HDR, that is, a statistical region containing those sample values of relatively high density (see Definition 1). We denote by $X_{ij}\ (x_{ij})$ the $j$-th component of $X_{i}\ (x_{i})$, $j=1,\dots,d$ and $i=1,\dots,n$. We restrict our discussion to continuous random variables $X$ and, unless otherwise stated, we denote by $f$ their probability density function (PDF) and by $F$ their cumulative density function (CDF).
Definition 1 (Highest-density region; Hyndman, (1996))
@@
One of the most distinctive properties of HDRs is that, among all regions of probability coverage $100(1-\alpha)\%$, the HDR has the smallest possible volume. The notion of “smallest” is to be understood with respect to some measure such as the usual Lebesgue measure; in the continuous one-dimensional case that would lead to the shortest-length set, while in two dimensions that would be the smallest-area set. It also follows from the definition that the boundary of an HDR consists of those values of the sample space with equal density. Hence a plot of a bivariate HDR has as the boundary a contour plot (see Figure 2).
@@ 2.1 Density quantile approach
The study of HDRs has been largely enhanced by Hyndman, who proposed the density quantile approach (outlined in Proposition 1) to compute multivariate HDRs (Hyndman,, 1996). Today, this still represents the typical strategy and involves estimating the density.
Proposition 1 (Hyndman, (1996))
Let $\{f(x_{1}),\dots,f(x_{m})\}$ be a sample of independent observations of size $m$ of the random variable $Y=f(X)$, with $f$ a bounded and continuous function in $x$. Consider the ordered sample $\{f_{(1)},\dots,f_{(m)}\}$ with $f_{(j)}$ the $j$-th largest among the $f(x_{i})$’s so that $f_{(j)}$ is the $(j/m)$ sample quantile of $Y$. Then, given a constant $\alpha\in[0,1]$, and denoted with $\lfloor j\rfloor$ the greatest integer less than or equal to $j$, we have that, in probability,
[…]
Basically, the HDR is derived based on the sample quantile of the density $f$. However, the density function itself is often unknown and we have to estimate it based on a set of available iid observations $s_{n}\doteq\{x_{1},\dots,x_{n}\}$. In this case, the $100(1-\alpha)\%$ HDR can be estimated as
$\hat{C}_{n}(\hat{f}_{\alpha})$ $\doteq C(x\!:f_{n}(x)>f_{\lfloor\alpha n\rfloor}),$ (1)
@@ 2.2 One-class neighbor machines (OCNM) approach
An alternative approach to HDR estimation, inspired by the theory of support vector machines (Schölkopf et al.,, 2001) and the notion of neighborhood measures, has been introduced in the machine learning literature by Munoz and Moguerza, (2006). The procedure is outlined in Proposition 2, after introducing neighborhood measures in Definition 2.
Definition 2 (Neighbourhood Measure; Munoz and Moguerza, (2006))
@@
Proposition 2 (Munoz and Moguerza, (2006))
@@
[…]
@@
[…]
where $C(f_{\nu})$ is the minimum-volume set such that $\mathbb{P}(C(f_{\nu}))\geq\nu$.
@@ 2.3 Neighborhood-quantile approach
Theorem 1
Let $X$ be a continuous random variable defined on $\mathbb{R}^{d}$ with density function $f$, and consider a set $s_{n}=\{x_{1},\dots,x_{n}\}\in S_{n}$ of size $n$ (drawn from $f$). Assume $g\!:S_{n}\times\mathbb{R}^{d}\to\mathbb{R}$ is a neighbourhood measure. Then an estimate of the $100(1-\alpha)\%$ HDR can be obtained as:
[…]
where $g(x_{\lfloor\alpha n\rfloor},s_{n})$ is the $\alpha$-quantile of the sample $\{g(x_{1},s_{n}),\dots,g(x_{n},s_{n})\}$, and $\lfloor j\rfloor$ denotes the greatest integer less than or equal to $j$.
The proof is straightforward when considering the relationship between the region $C^{\text{OCNM}}_{n}$ as defined in Proposition 2 and the region $C_{n}$ as defined in Theorem 1. In fact, without loss of generality, taken $g$ to be a sparsity measure, and noticing that $\nu$ plays the role of $1-\alpha$, we can observe that:
@@
=$=\{x\!:\operatorname{sign}(g(x_{\lceil\nu n\rceil},s_{n})-g(x,s_{n}))\geq 0\}$
=$=\{x\!:g(x,s_{n})\leq g(x_{\lceil(1-\alpha)n\rceil},s_{n})\}$
@@
Remark If $g$ is chosen to be a concentration measure, then to ensure a $100(1-\alpha)\%$ coverage (notice that in this case $\mathbb{P}(C(f_{\nu}))=1-\nu$), the decision value $\rho^{*}$ induced by the concentration measure is given by $\rho^{*}=g(x_{\lfloor\alpha n\rfloor},s_{n})$.
Noticing that the density represents a concentration measure, the resemblance with the density quantile approach in Proposition 1 should now be clear. In fact, from the perspective of the density quantile approach, we can view Theorem 1 as a generalization of Proposition 1, where $f$ is replaced by any function $g$ that satisfies the criteria of neighborhood measures. Intuitively, provided that the density function $f$ is replaced by a function that preserves the order induced in the sample by $f$, the estimated HDR is asymptotically valid. Neighborhood measures ensure this ranking. To see it, without loss of generality, consider a sparsity measure $g$ and a sample $s_{n}=\{x_{1},\dots,x_{(1-\alpha)n},\dots,x_{n}\}$ ordered so that
@@
Remark Proposition 1 represents a special case of Theorem 1 with $g(x,s_{n})$ being either: (a) a concentration measure $g(x,s_{n})\propto\hat{f}(x,s_{n})$ or; (b) a sparsity measure $g(x,s_{n})\propto\frac{1}{\hat{f}(x,s_{n})}$, where $\hat{f}$ can be any consistent density estimator. Among the plethora of density estimators, in this work we focus on KDE (Parzen,, 1962), and use it as a benchmark measure for the proposed comparators in Section 3. Specifically, given a set of iid observations $s_{n}\doteq\{x_{1},\dots,x_{n}\}$ drawn from an unknown target density $f$, the KDE measure $M_{0}(x,s_{n},h)$ at the location $x$ is defined as
$M_{0}(x,s_{n},h)=\frac{1}{n}\sum_{i=1}^{n}K_{h}\left(\lVert x-x_{i}\rVert\right)=\frac{1}{nh^{d}}\sum_{i=1}^{n}K\left(\frac{\lVert x-x_{i}\rVert}{h}\right),$ (2)
@@
3 Alternative Measures for Building HDRs
We now discuss a number of neighborhood measures that could be used to estimate HDRs. We repurpose some measures that have been successfully employed in areas such as classification or clustering, and introduce some novel ideas, included a copula-based approach.

### v2

# Alternative Approaches for Estimating Highest-Density Regions

@@ Abstract
Among the variety of statistical intervals, highest-density regions (HDRs) stand out for their ability to effectively summarize a distribution or sample, unveiling its distinctive and salient features. An HDR represents the minimum size set that satisfies a certain probability coverage, and current methods for their computation require knowledge or estimation of the underlying probability distribution or density $f$. In this work, we illustrate a broader framework for computing HDRs, which generalizes the classical density quantile method introduced in the seminal paper of Hyndman, (1996). The framework is based on neighbourhood measures, i.e., measures that preserve the order induced in the sample by $f$, and include the density $f$ as a special case. We explore a number of suitable distance-based measures, such as the $k$-nearest neighborhood distance, and some probabilistic variants based on copula models. An extensive comparison is provided, showing the advantages of the copula-based strategy, especially in those scenarios that exhibit complex structures (e.g., multimodalities or particular dependencies). Finally, we discuss the practical implications of our findings for estimating HDRs in real-world applications.
Keywords: Anomaly detection, Copula models, Density estimation, $k$-nearest neighborhood, Statistical intervals
@@ 1 Introduction
A ubiquitous problem in statistics is to derive statistical intervals or regions–especially in the multivariate setting–for population parameters or other unknown quantities. Their role is to provide a way to quantify and describe the uncertainty about a quantity of interest, or simply a way to summarize the information contained in a distribution. Statistical regions may address different problems. For example, a confidence interval (CI) describes the uncertainty related to an estimate for an unknown parameter, while a prediction interval provides bounds for one or more future observations. Alternatively, a tolerance interval is the interval expected to contain a specified proportion of the sampled population. In a Bayesian setting, highest posterior density or credible regions provide, in a natural way, set estimates for a specific parameter (Box and Tiao,, 1992; Turkkan and Pham-Gia,, 1993). We refer to Meeker et al., (2017) and Krishnamoorthy and Mathew, (2009) for an overview. Furthermore, even when the focus is on a specific type of interval, e.g., a two-sided 95% prediction interval, questions on how this should be defined may still arise. Should we use the interval symmetric about the mean or the interval of shortest length, among others? Although each of these intervals has 95% coverage, they may all be different. Consider, for example, the case of nonsymmetric and/or multimodal distributions, with an illustration given in Figure 1. An additional layer of complexity arises in multivariate settings, where there are no unique agreed definitions, and generalizations include the concept of simultaneous CIs (Guilbaud,, 2008), multivariate CIs (Korpela et al.,, 2017), or different definitions for multivariate quantiles (Figalli,, 2018; Cai,, 2010; Coblenz et al.,, 2018), among others.
@@
The multivariate setting will be the target of this work, with a focus on bivariate distributions. In particular, our interest is devoted to statistical regions for summarizing probability distributions in the form of highest-density regions (HDRs; Hyndman,, 1996). Statistical regions other than HDRs are beyond the scope of the present work, and we refer to Meeker et al., (2017) and  Krishnamoorthy and Mathew, (2009) for a comprehensive survey on the broader topic. As the name suggests, an HDR specifies the set of points of highest density: the density for points inside the region must be higher than that for points outside it. More specifically, considering a $d$-dimensional continuous variable of interest $X\in\mathbb{R}^{d},d\geq 1$, with probability density function $f$, the problem is to estimate minimum volume sets of the form $C(f_{\alpha})=\{x\!:f(x)\geq f_{\alpha}\}$, such that $P(X\in C(f_{\alpha}))\geq 1-\alpha$, where $1-\alpha$, with $\alpha\in(0,1)$, represents a prespecified coverage probability. Although they share substantial similarities with multivariate quantiles, estimating an HDR differs from estimating level sets in that one is interested in specifying a probability content rather than the level directly. This complicates the problem, and we refer to Doss and Weng, (2018) for more details.
@@
To detect abnormal observations from a sample: if a data point does not belong to a region of “normal” data (the HDR), then it is regarded anomalous (see e.g., Steinwart et al.,, 2005, and references therein).
@@
This work will primarily be driven by the problem of anomaly detection, characterizing a broad spectrum of applied domains, going from astrophysics to diagnostics and sport analytics. More specifically, our interest is to develop a valid and efficient framework in support of the worldwide doping detection mission headed by the World Anti-Doping Agency (WADA,, 2021). In practice, WADA’s current analytical implementation is based on identifying reference ranges that discriminate well between normal and abnormal values for predefined biomarkers of interest (Sottas et al.,, 2007). This is done following a univariate approach, with reference ranges, in the form of equal-tailed intervals, derived for each biomarker separately. Clearly, addressing this problem over increased dimensions presents significant challenges, including the presence of data with complex dependence structures, in addition to distributional multimodalities or skewness.
Due to their flexibility “to convey both multimodality and asymmetry”, HDRs are argued to be a more effective summary of the distribution (Hyndman,, 1995). In the case of unimodal symmetric distributions, such as the normal distribution, an HDR coincides with the usual probability region symmetric about the mean, spanning the $\alpha/2$ and $21-\alpha/$ 2 quantiles. However, in the case of a multimodal distribution, it may consist of several disjoint subregions, each containing a local mode. This provides useful information that could not be traced by other probability regions such as an equal-tailed interval (see Figure 1 for an illustrative example).
@@
As the name suggests, estimating an HDR is based on knowing the density function $f$ of the variable of interest $X$. However, as typically occurs in practice, this quantity is unknown and the estimation of an HDR requires estimating $f$ first. The seminal paper of Hyndman, (1996) discusses the density-quantile approach for computing HDRs in such settings. Although for unidimensional problems this task can be achieved very accurately using methods such as the kernel density estimator (KDE; Parzen,, 1962) or the local likelihood approach (Hjort and Jones,, 1996), it may be inefficient for multidimensional problems (Liu et al.,, 2007). In fact, over increased dimensions, KDE suffers from the difficulty of finding optimal kernel functions and the corresponding bandwidths (i.e., the smoothing parameters). In particular, bandwidth selection in KDE is recognized as the most crucial and difficult step (see e.g., Chapter 2 in Wand and Jones, 1994a, ), with no definite and widely accepted solution. Furthermore, high-dimensional data also pose challenges from an algorithmic/computational perspective when deriving the associated HDR.
@@
In this work, we illustrate a broad framework for estimating HDRs, which generalizes the current density-quantile approach implementable on the basis of a consistent estimator of the (multivariate) density (Hyndman,, 1996). The proposed framework is based on neighbourhood measures (Munoz and Moguerza,, 2006), that is, measures that preserve the order induced in the sample by the density function. Notably, it includes the widely-used density estimation procedure as a special case. We then elaborate on and evaluate a number of suitable probabilistic- and distance-based measures, including a variation of the measure adopted by the k-nearest neighbors algorithm. In particular, motivated by the ubiquitous role of copula modeling (Nelsen,, 2006) in modern statistics, among probabilistic-based measures, we explore the use of copulae in an HDR estimation context. Interestingly, copulae introduce more flexibility to deal with multivariate random vectors, by separately estimating the marginals and their dependence structure, i.e., the copula model. In addition, by placing a strong focus on the dependence model, a copula approach has the advantage of better capturing data specificities, especially when these exhibit complex relationships such as asymmetric and/or tail dependencies.
The remainder of this manuscript is organized as follows. In Section 2, we introduce the problem of interest and review existing methods for HDR estimation. The general neighborhood-quantile framework is described in Section 2.3. In Section 3, we discuss and propose alternative measures in the context of HDRs, including some variants based on copulae. Section 4 provides a comprehensive comparison among the introduced measures and the standard kernel density estimator. Empirical studies are focused on bivariate scenarios that vary according to the complexity of the data (marginal and dependence structure, multimodality, etc) and the sample size. A distribution on a compact support, namely, the Dirichlet distribution, is also considered to support practical applications with compositional data. An application to the MAGIC data set, which classifies high-energy Gamma particles in the atmosphere (Bock et al.,, 2004), is illustrated in Section 4.5. We conclude in Section 5 by summarizing the main findings and discussing their implications for estimating HDRs in real-world problems.
2 A general framework for HDR estimation
@@
Let $\{X_{1},\dots,X_{n}\}$ be a sample of $n$ independent and identically distributed (iid) replications of a random variable $X$ defined on $\mathbb{R}^{d}$, with $d\geq 1$. In our problem, we assume to have access to a sample $s_{n}=\{x_{1},\dots,x_{n}\}\in S_{n}$ of their actual realizations, with $S_{n}$ the sample space. We then wish to use $s_{n}$ for estimating an HDR, that is, a statistical region containing those sample values of relatively high density (see Definition 1, attributed to Hyndman,, 1996). We denote by $X_{ij}\ (x_{ij})$ the $j$-th component of $X_{i}\ (x_{i})$, for $j=1,\dots,d$ and $i=1,\dots,n$. We restrict our discussion to continuous random variables $X$ and, unless otherwise stated, we denote by $f$ their probability density function (PDF) and by $F$ their cumulative density function (CDF).
Definition 1 (Highest-density region; Hyndman,, 1996).
@@
One of the most distinctive properties of HDRs is that, among all regions of probability coverage $100(1-\alpha)\%$, the HDR has the smallest possible volume. The notion of “smallest” is to be understood with respect to some measure such as the usual Lebesgue measure; in the continuous one-dimensional case that would lead to the shortest-length set, while in two dimensions that would be the smallest-area set. It also follows from the definition that the boundary of an HDR consists of those values of the sample space with equal density. Hence a plot of a bivariate HDR has as the boundary a contour plot.
@@ 2.1 Density quantile approach
The study of HDRs has been largely enhanced by Hyndman, who proposed the density quantile approach (outlined in Proposition 1) to estimate multivariate HDRs (Hyndman,, 1996). Today, this still represents the typical strategy and involves estimating the density.
Proposition 1 (Hyndman,, 1996).
Let $\{f(x_{1}),\dots,f(x_{m})\}$ be a sample of independent observations of size $m$ of the random variable $Y=f(X)$, with $f$ a bounded and continuous function in $x$. Consider the ordered sample $\{f_{(1)},\dots,f_{(m)}\}$ with $f_{(j)}$ the $j$-th largest among the $f(x_{i})$’s so that $f_{(j)}$ is the $(j/m)$ sample quantile of $Y$. Then, given a constant $\alpha\in[0,1]$, and denoted with $\lfloor j\rfloor$ the greatest integer less than or equal to $j$, in probability,
$\hat{f}_{\alpha}$ $\doteq f_{(\lfloor\alpha m\rfloor)}\to f_{\alpha}\quad\text{as}\ m\to\infty,$
$C_{m}(\hat{f}_{\alpha})$$\doteq\{x\!:f(x)>\hat{f}_{\alpha}\}\to C(f_{\alpha})\quad\text{as}\ m\to\infty.$
Basically, the HDR is derived based on the sample quantile of the density $f$. However, the density function itself is often unknown and one has to estimate it based on a set of available iid observations $s_{n}\doteq\{x_{1},\dots,x_{n}\}$. In this case, the $100(1-\alpha)\%$ HDR can be estimated as
$\hat{C}_{n}(\hat{f}_{\alpha})$$\doteq\{x\!:f_{n}(x)>f_{(\lfloor\alpha n\rfloor)}\},$ (1)
@@ 2.2 One-class neighbor machines (OCNM) approach
An alternative approach to HDR estimation, inspired by the theory of support vector machines (Schölkopf et al.,, 2001) has been introduced in the machine learning literature by Munoz and Moguerza, (2006). The procedure is outlined in Proposition 2, and involves the notion of neighborhood measures (see Definition 2, attributed to Munoz and Moguerza,, 2006).
Definition 2 (Neighbourhood measure; Munoz and Moguerza,, 2006).
@@
Proposition 2 (Munoz and Moguerza,, 2006).
@@
$\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(h(x_{i})\neq-1)=\nu\right)\to 1\quad\text{as}\ n\to\infty.$
@@
$C^{\text{OCNM}}_{n}\doteq\{x\!:h(x)\geq 0\}\to C(f_{\nu})\doteq\{x\!:f(x)\geq f_{\nu}\}\quad\text{as}\ n\to\infty,$
where $C(f_{\nu})$ is the minimum-volume set such that $\mathbb{P}(C(f_{\nu}))\geq\nu$, with $\nu\in[0,1]$.
@@ 2.3 Neighborhood-quantile approach
Theorem 1.
Let $X$ be a continuous random variable defined on $\mathbb{R}^{d}$ with density function $f$, and consider a set $s_{n}=\{x_{1},\dots,x_{n}\}\in S_{n}$ of size $n$ (drawn from $f$). Assume $g\!:\mathbb{R}^{d}\times S_{n}\to\mathbb{R}$ is a neighbourhood measure. Then an estimate of the $100(1-\alpha)\%$ HDR can be obtained as:
$C_{n}$ =$=\{x\!:g(x,s_{n})\leq g(x_{(\lceil(1-\alpha)n\rceil)},s_{n})\}\quad\text{if}\ g\ \text{is a sparsity measure},$
$C_{n}$ =$=\{x\!:g(x,s_{n})\geq g(x_{(\lfloor\alpha n\rfloor)},s_{n})\}\quad\quad\ \ \text{if}\ g\ \text{is a concentration measure},$
where $g(x_{(\lfloor\alpha n\rfloor)},s_{n})$ is the $\alpha$-quantile of the sample $\{g(x_{1},s_{n}),\dots,g(x_{n},s_{n})\}$, and $\lfloor j\rfloor$ denotes the greatest integer less than or equal to $j$.
The proof is straightforward when considering the relationship between the region $C^{\text{OCNM}}_{n}$ as defined in Proposition 2 and the region $C_{n}$ as defined in Theorem 1. In fact, without loss of generality, taking $g$ to be a sparsity measure, and noticing that $\nu$ plays the role of $1-\alpha$, one can observe that:
@@
=$=\{x\!:\operatorname{sign}(g(x_{(\lceil\nu n\rceil)},s_{n})-g(x,s_{n}))\geq 0\}$
=$=\{x\!:g(x,s_{n})\leq g(x_{(\lceil(1-\alpha)n\rceil)},s_{n})\}$
@@
Remark If $g$ is chosen to be a concentration measure, then to ensure a $100(1-\alpha)\%$ coverage (notice that in this case $\mathbb{P}(C(f_{\nu}))=1-\nu$), the decision value $\rho^{*}$ induced by the concentration measure is given by $\rho^{*}=g(x_{(\lfloor\alpha n\rfloor)},s_{n})$.
Noticing that the density represents a concentration measure, the resemblance with the density quantile approach in Proposition 1 should now be clear. In fact, from the perspective of the density quantile approach, one can view Theorem 1 as a generalization of Proposition 1, where $f$ is replaced by any function $g$ that satisfies the criteria of neighborhood measures. Intuitively, provided that the density function $f$ is replaced by a function that preserves the order induced in the sample by $f$, the estimated HDR is asymptotically valid. Neighborhood measures ensure this ranking. To see it, without loss of generality, consider a sparsity measure $g$ and a sample $s_{n}=\{x_{1},\dots,x_{(1-\alpha)n},\dots,x_{n}\}$ ordered so that
@@
Remark Proposition 1 is a special case of Theorem 1 with $g(x,s_{n})$ being either: (a) a concentration measure $g(x,s_{n})\propto\hat{f}(x,s_{n})$ or; (b) a sparsity measure $g(x,s_{n})\propto\frac{1}{\hat{f}(x,s_{n})}$, where $\hat{f}$ can be any consistent density estimator. Among the plethora of density estimators, in this work we focus on KDE (Parzen,, 1962), and use it as a benchmark measure for the proposed comparators in Section 3. Specifically, given a set of iid observations $s_{n}\doteq\{x_{1},\dots,x_{n}\}$ drawn from an unknown target density $f$, the KDE measure $M_{0}(x,s_{n},h)$ at the location $x$ is defined as
$M_{0}(x,s_{n},h)=\frac{1}{nh^{d}}\sum_{i=1}^{n}K\left(\frac{\lVert x-x_{i}\rVert}{h}\right),$ (2)
@@
3 Alternative measures for estimating HDRs
We now propose a number of neighborhood measures that could be used to estimate HDRs. We elaborate on some measures that have been successfully employed in areas such as classification or clustering and introduce some novel ideas, including a copula-based approach. Although some of the discussed distances are popular in existing statistical domains, e.g., the $k$-nearest neighbors distance and its use in classification and regression, these have not been considered in the context of HDRs.